# Training Hybrid Deep Quantum Neural Networks for Efficient Reinforcement Learning

### v5

@@ C.2 Detailed Graphs with seeds
Figure 5: Moving-average ($100$-episode window) return curves for four PQC configurations. Each colored trace corresponds to one random seed, consistent across the experiments.
Appendix D A word on the technical stack: NVIDIA CUDA-Q
We adopted CUDA-Q [^6] https://developer.nvidia.com/cuda-quantum as our quantum-simulation engine for one main reason: the GPU is a light weight and agile way to develop and carry out small-scale tests for new model architectures with a small number of qubits involved. Fine-grained quantitative investigations could be carried out easily in this way to understand the behaviors of novel hybrid quantum-classical machine learning architectures by limiting the qubit count, offering direct insights into their foundation of learning and information processing flows. CUDA-Q provides GPU-accelerated tensor-network and state-vector simulators for quick proof-of-concept demonstrations on GPUs. The hDQNN architectures validated in through this could be straightforwardly scaled towards beyond classical large models (with a medium or large qubit count) running on hybrid systems having classical computing resources (CPU/GPU) interconnected with physical (noisy) QPUs.
D.1 Hardware and system software.
@@ Appendix D A word on the technical stack: NVIDIA CUDA-Q
All experiments were executed independently on a dedicated NVIDIA A100-SXM4 GPU (40 GB HBM2e). The software stack comprised:
@@
D.2 Supplementary figures

### v6

@@ E.3 Detailed Graphs with seeds
Figure 6: Moving‐average (100100‐episode window) return curves for four PQC configurations. Each colored trace corresponds to one random seed, consistent across the experiments.
E.4 Other supplementary figures
Appendix F Technical stack
F.1 Hardware and system software.
@@ Appendix F Technical stack
All numerical experiments were executed independently on a dedicated NVIDIA A100-SXM4 GPU (40 GB HBM2e). The software stack comprised:
@@
F.2 NVIDIA CUDA-Q
We adopted CUDA-Q [^7] https://developer.nvidia.com/cuda-quantum as our quantum-simulation engine for one main reason: the GPU is a light weight and agile way to develop and carry out small-scale tests for new model architectures with a small number of qubits involved. Fine-grained quantitative investigations could be carried out easily in this way to understand the behaviours of novel hybrid quantum-classical machine learning architectures by limiting the qubit count, offering direct insights into their foundation of learning and information processing flows. CUDA-Q provides GPU-accelerated tensor-network and state-vector simulators for quick proof-of-concept demonstrations on GPUs. The hDQNN architectures validated in through this could be straightforwardly scaled towards beyond classical large models (with a medium or large qubit count) running on hybrid systems having classical computing resources (CPU/GPU) interconnected with physical (noisy) QPUs.
F.3 Resource footprint of the surrogate
A 10-qubit qtDNN has 473,098473{,}098 trainable parameters ($\approx 1.80$ MB of FP32 weights). Adam adds two moments of equal size ($\approx 3.60$ MB), keeping the persistent footprint under $6$ MB. On our A100 40GB runs, qtDNN updates account for ${\sim}2\%$ of wall-clock (median across seeds).
F.4 Deployment regime
All experiments execute the PQC with CUDA‑Q’s shot‑based simulator augmented by bit/phase‑flip noise channels to emulate NISQ behaviour. At test time the policy executes the same noisy PQC (on calibrated hardware or its simulator). The qtDNN is never used for inference.
F.5 Noise sources in NISQ-compatible training
Throughout training, the system operates under three distinct noise regimes:
  • Sampling noise, arising from finite-shot PQC measurements, which produce stochastic bit-strings even in the absence of gate errors (we considered $100$ and $1000$ shots in the ablation study).
  • Gate noise, stemming from hardware-level imperfections and is modeled via bit-flip and phase-flip channels with realistic error rates ($\sim 10^{-3}$ in the ablation study).
  • Exploration noise, entering through stochastic policy perturbations used during environment interaction via application of Gaussian noise.
The qtDNN surrogate is trained directly on the resulting noisy bit-strings and thus inherits robustness to all three sources during gradient-based optimisation.
Appendix G Quantum advantage of PQC over classical architectures
G.1 Compact multi‑frequency structure.
Expectation values of data‑encoded PQCs admit a finite Fourier series; the largest frequency scales at most linearly with circuit depth. A shallow ReLU layer can match the same spectra only by increasing width rapidly with frequency and dimension, implying an exponential width gap in the multi‑dimensional setting (cf. Schuld et al. [2021], Montúfar et al. [2014]).
G.2 Entanglement as an inductive bias.
The entangling feature map induces a data‑dependent quantum kernel that captures non‑local correlations which comparable‑size random‑feature maps struggle to reproduce; when the geometric discrepancy is large, one can construct datasets on which the quantum model achieves strictly lower prediction error (cf. Huang et al. [2021b]).
Empirically, replacing the PQC with an FC layer of identical width lowers mean return by +296+296 at 4k episodes and +148+148 at 7.2k episodes (Table 2), consistent with these lenses.